# Learning Value of Information towards Joint Communication and Control in 6G V2X

### v2

@@ VI-B3 Method C
We estimate EVoMI and IVoMI for the vehicle-following control problem as given in Case 6. The velocity trajectory of the preceding vehicle $i-1$ is extracted from the Next Generation Simulation (NGSIM) dataset [40] when training the following vehicle. The optimal policy $\tilde{\pi}_{i}^{*}$ for the augmented-state SSDP in Case 4 is learned by the Twin Delayed Deep Deterministic policy gradient (TD3) algorithm [41]. This policy is also used as the acting policy $\pi_{i}$ when the predecessor’s acceleration is not available. In this case, $acc_{i-1,k}$ is set to 00 in the state.
The EVoMI is $-3.12$, which is calculated by evaluating the policy over $100$ test episodes in which $acc_{i-1,k}$ is either consistently unavailable or consistently available throughout the entire episode, and then computing the difference in the average total reward between the two settings.
@@ VII-B3 Across Multiple Control Intervals - When and What to Communicate
The “When to Communicate” and “What to Communicate” decisions can also be made statically only once at control interval 00 or semi-statically for every $C>1$ control intervals. In this case, the subset of information or processed information will always be transmitted at every control interval within the effective time range of the decision. An example is choosing the information topology for a platoon [35].
@@
We consider a typical urban C-V2X network, where V2V links coexist with V2I links. A V2I link connects a vehicle to the BS and can be used for high-throughput services. A V2V link connects a pair of preceding vehicle $i-1$ and following vehicle $i$ for periodic transmission of cooperative awareness messages (CAM) to support advanced driving services. Specifically, at the beginning of each control interval $k$, the acceleration $acc_{i-1,k}$ of the predecessor $i-1$ is sampled to form the CAM and buffered in a queue before transmission to the corresponding follower $i$.
@@
At each communication interval $(k,t)$, the predecessor $i-1$ transmits the CAM generated at the beginning of each control interval $k$. Any CAM that is not fully transmitted by the end of the last communication interval $(k,T-1)$ of control interval $k$ will be discarded. In other words, any CAM not fully delivered during the previous control interval $k-1$ is replaced by newly generated CAM at the beginning of control interval $k$.
@@
As introduced in Case 1, the control agent in follower $i$ determines its control input based on the local sensory information $ep_{i,k}$, $ev_{i,k}$, $acc_{i,k}$, as well as the remote sensory information received from its predecessor $i-1$.
@@
There is a communication agent in each predecessor $i-1$ to make communication decisions, which fall into one or more of the following categories:
@@
  • Every Communication Interval - How to Communicate: The communication agent makes RRM decisions at each communication interval $(k,t)$. To enhance spectrum utilization, we consider a setting where $L$ V2V links share the uplink resources orthogonally allocated to $M$ V2I links. With a slight abuse of notation, each V2V link, connecting a predecessor $i-1$ and its corresponding follower $i$, is indexed by $i-1\in\mathcal{L}=\{1,3,5,\ldots,2L-1\}$. Without loss of generality, we assume that every V2I link $m\in\mathcal{M}=\{0,\cdots,M-1\}$ is pre-assigned sub-channel $m$ and transmits with a constant power $P^{\rm I}_{m}$. Moreover, one or more V2V links can reuse the sub-channels of the V2I links for CAM transmission. We use the binary allocation indicator $\theta_{i-1,m,(k,t)}\in\{0,1\}$ to indicate whether V2V link $i-1$ occupies sub-channel $m$ at communication interval $(k,t)$ or not. Moreover, we consider that each V2V link $i-1$ occupies at most one sub-channel, i.e., $\sum_{m=0}^{M-1}\theta_{i-1,m,(k,t)}\leq 1$. Let $P^{\rm V}_{i-1,m,(k,t)}$ denote the transmit power of V2V link $i-1$ over the sub-channel $m$ at communication interval $(k,t)$. Therefore, each communication agent $i-1$ has to make the subchannel allocation $\{\theta_{i-1,m,(k,t)}\}_{m\in\mathcal{M}}$ and power allocation $\{P^{\rm V}_{i-1,m,(k,t)}\}_{m\in\mathcal{M}}$ decisions at each communication interval $(k,t)$.
@@
  • Every Control Interval - When to Communicate: Let $\phi_{i-1,k}\in\{0,1\}$ be the binary transmission indicator to indicate whether the sampled CAM $acc_{i-1,k}$ will be transmitted to the follower $i$ or not. At each control interval $k$, each communication agent $i-1$ can make decisions on $\phi_{i-1,k}$.
@@
  • Across Multiple Control Intervals - When to Communicate: Each communication agent $i-1$ makes a one-time decision on whether it will transmit its acceleration or not to its follower $i$. Once the decision is made, the predecessor will always transmit $acc_{i-1,k}$ or not at each control interval $k$.
@@ VII-C Optimization Objective
The instantaneous channel gain of V2V link $i-1$ over sub-channel $m$ (occupied by V2I link $m$) at communication interval $(k,t)$ is denoted by $G_{i-1,m,(k,t)}$. $G_{i-1,m,(k,t)}$ remains constant within a communication interval, and can be written as
@@
Similarly, let $G_{m,(k,t)}$ denote the instantaneous channel gain of V2I link $m$; $G_{i-1,B,m,(k,t)}$ the interference channel gain from V2V link $i-1$ transmitter at predecessor $i-1$ to V2I link $m$ receiver at the BS; $G_{B,i-1,m,(k,t)}$ the interference channel gain from V2I link $m$ transmitter at a vehicle to V2V link $i-1$ receiver at follower $i$; and $G_{j-1,i-1,m,(k,t)}$ the interference channel gain from V2V link $j-1$ transmitter at predecessor $j-1$ to V2V link $i-1$ receiver at follower $i$ over the sub-channel $m$.
The SINR $\gamma_{m,(k,t)}$ of V2I link $m$ and the SINR $\gamma_{i-1,m,(k,t)}$ of V2V link $i-1$ on sub-channel $m$ at communication interval $(k,t)$ are derived by
@@
respectively. $\sigma^{2}$ is the power of channel noise which satisfies the independent Gaussian distribution with a zero mean value. $I_{i-1,m,(k,t)}$ in (38) is the total interference power received by V2V link $i-1$ over sub-channel $m$, where
@@
$\sum\limits_{j-1\in\mathcal{L}\backslash\{i-1\}}\phi_{i-1,k}\theta_{j-1,m,(k,t)}P^{\rm V}_{j-1,m,(k,t)}G_{j-1,i-1,m,(k,t)}.$
The instantaneous data rates $r_{m,(k,t)}$ and $r_{i-1,(k,t)}$ of V2I link $m$ and V2V link $i-1$ at communication interval $(k,t)$ are respectively derived as
$r_{m,(k,t)}=B\log_{2}(1+\gamma_{m,(k,t)}),$ (39)
@@
Let $r^{\rm CAM}_{i-1,(k,t)}$ denote the transmission rate of V2V link $i-1$ in terms of CAM at communication interval $(k,t)$, which is given by
$r^{\rm CAM}_{i-1,(k,t)}=\frac{r_{i-1,(k,t)}}{N_{c}},$ (41)
@@
Each predecessor $i-1\in\mathcal{L}$ maintains a buffer to store its CAMs. Let $q^{\rm CAM}_{i-1,(k,t)}$ denote the queue length, i.e., the number of CAMs in the buffer of predecessor $i-1$ at communication interval $(k,t)$. At the beginning of the first communication interval $(k,0)$ of each control interval $k\in\mathcal{K}$, predecessor $i-1$ samples the driving status for control interval $k$ and generates a CAM. If the communication agent decides to transmit this CAM ($\phi_{i-1,k}=1$), it is stored in the buffer. During each communication interval $(k,t)$, where $t\in\mathcal{T}$, the queue length $q^{\rm CAM}_{i-1,(k,t)}$ is decreased by $\Delta t\times r^{\rm CAM}_{i-1,(k,t)}$, which represents the number of CAMs transmitted during this communication interval.
Since any CAM not fully transmitted by the end of the last communication interval $(k,T-1)$ is discarded, the queue is reset at the start of each new control interval $k+1$. Specifically, at communication interval $(k+1,0)$, $q^{\rm CAM}_{i-1,(k+1,0)}$ is set to $1$ if a CAM is scheduled for transmission ($\phi_{i-1,k+1}=1$), or 00 otherwise ($\phi_{i-1,k+1}=0$). Because the queue length changes abruptly at the boundaries between control intervals, we use $q^{\rm CAM}_{i-1,(k,T)}$ to denote the queue length at the end of control interval $k$, i.e., after the final communication interval $(k,T-1)$. Note that $q^{\rm CAM}_{i-1,(k,T)}\neq q^{\rm CAM}_{i,(k+1,0)}$ due to this discontinuity.
@@
$\left\{\begin{array}[]{ll}\phi_{i-1,k},&\mathrm{if}\quad t=0\\ \max\left[0,q^{\rm CAM}_{i-1,(k,t-1)}-r^{\rm CAM}_{i-1,(k,t-1)}\Delta t\right],&\mathrm{if}\quad 0<t\leq T\\ \end{array}\right..$ (44)
@@
With option (2), when the most recent available CAM of follower $i$ at the beginning of control interval $k$ corresponds to the CAM sampled at the predecessor $i-1$ in control interval $k-\tau_{i,k}$, we refer to $\tau_{i,k}$ as the observation delay at V2V receiver $i$, which is derived from the queue length as
@@
$J^{\mathrm{CM}}=\kappa_{1}\sum_{m=0}^{M-1}\mathrm{E}_{\pi^{\rm CM}}\left[\sum_{k=0}^{K-1}\sum_{t=0}^{T-1}\gamma^{kT+t}r_{m,(k,t)}\right]+\kappa_{2}\sum_{i-1\in\mathcal{L}}\Xi_{i},$ (48)
@@ VII-D VoI-oriented Optimization for Vehicular Communication
    $r^{\rm CM}_{(k,t)}=\kappa_{1}\left(\sum_{t=0}^{T-1}\gamma^{t}r_{(k,t)}^{\mathrm{CM\_trad}}\right)+\kappa_{2}\xi_{k+1},$ (50)
@@
Note that the $r_{(k,t)}^{\mathrm{CM\_trad}}$ component is present in the reward of every communication interval, while the IVoII component, i.e., $\xi_{k+1}$, is only present in the last communication interval of each control interval when $t=T-1$. The imperfect observation received by the control agent at control interval $k+1$ is affected by the degree of communication nonideality the transmitted data has experienced, which is in part a result of the sequence of decisions on communication actions $a_{(k,0)}^{\text{CM}},a_{(k,1)}^{\text{CM}},\ldots,a_{(k,T-1)}^{\text{CM}}$ that are selected during control interval $k$ by the communication agent.
@@
      $r^{\rm CM}_{(k,t)}$ $=\kappa_{1}\left(\sum_{m=0}^{M-1}\sum_{t=0}^{T-1}\gamma^{t}r_{m,(k,t)}\right)$
@@
      $\left\{\begin{array}[]{ll}\kappa_{1}\sum_{m=0}^{M-1}r_{m,(k,t)},&\ 0\leq t<T-1\\ \kappa_{1}\sum_{m=0}^{M-1}r_{m,(k,t)}+&\\ \kappa_{2}\sum_{i-1\in\mathcal{L}}\xi_{i,k+1},&\ t=T-1\\ \end{array}\right..$ (58)
@@ VIII Conclusion and Future Work
$\mathrm{Pr}(f^{S}(S_{k},a_{k},f^{W}(S_{k},a_{k},\xi_{k}))|S_{k},a_{k})$
$=$ $\mathrm{Pr}(f^{S}(S_{k},a_{k},f^{W}(S_{k},a_{k},\xi_{k}))|S_{k},a_{k},\{S_{k^{\prime}},a_{k^{\prime}}\}_{k^{\prime}\leq k-1}),$ (60)
which follows from the system’s state transition function $S_{k+1}=f^{S}(S_{k},a_{k},W_{k})$ given in Definition 1. In addition, $W_{k}=f^{W}(\tilde{W}_{k})$ according to the exogenous information function in Definition 1, while $\tilde{W}_{k}\subseteq\{S_{k},a_{k},\xi_{k}\}$ as stated in Theorem 1.
@@
$\mathrm{Pr}(\xi_{k}|S_{k},a_{k})$
$=$ $\mathrm{Pr}(\xi_{k}|S_{k},a_{k},\{S_{k^{\prime}},a_{k^{\prime}}\}_{k^{\prime}\leq k-1})$
$=$ $\mathrm{Pr}(\xi_{k}|S_{k},a_{k},\{S_{k^{\prime}},a_{k^{\prime}},W_{k^{\prime}}\}_{k^{\prime}\leq k-1}),$
@@
Since $W_{k}=f^{W}(S_{k},a_{k},\xi_{k})$, $\xi_{k}$ is a set of random variables whose distribution is independent of $S_{k}$ and $a_{k}$ by the expression of the formula. Thus, from (-A), we have
$\mathrm{Pr}(\xi_{k})=\mathrm{Pr}(\xi_{k}|\{S_{k^{\prime}},a_{k^{\prime}},W_{k^{\prime}}\}_{k^{\prime}\leq k-1}),$ (62)
which shows that the distribution of $\xi_{k}$ does not depend on the past states, actions and exogenous information, i.e., $\{S_{k^{\prime}},a_{k^{\prime}},W_{k^{\prime}}\}_{k^{\prime}\leq k-1}$.
@@
$=\begin{pmatrix}f^{S}(S_{k},a_{k},W_{k})\\ f^{W}(f^{S}(S_{k},a_{k},W_{k}),\tilde{\pi}({f^{S}(S_{k},a_{k},W_{k})}),\xi_{k})\end{pmatrix}$
$=f^{\tilde{S}}(\tilde{S}_{k},a_{k},\xi_{k}),$ (65)
where the exogenous information $\xi_{k}$ is an independent random variable with given distribution.
@@
=$\stackrel{{\scriptstyle(b)}}{{=}}\max_{\pi_{k}(S_{k})}\Biggl{\{}\mathrm{E}_{W_{k}}\Biggl{[}r\bigl{(}\tilde{S}_{k},\pi_{k}(S_{k})\bigr{)}+\mathrm{E}_{W_{k+1}}\Bigl{[}\tilde{v}_{k+1}^{*}$
@@
=$\stackrel{{\scriptstyle(c)}}{{=}}\max_{\pi_{k}(S_{k})}\Biggl{\{}\mathrm{E}_{W_{k}}\Biggl{[}r\bigl{(}\tilde{S}_{k},\pi_{k}(S_{k})\bigr{)}+\mathrm{E}_{S_{k+1}}\Bigl{[}\mathrm{E}_{W_{k+1}}\Bigl{[}\tilde{v}_{k+1}^{*}$
@@
=$\stackrel{{\scriptstyle(d)}}{{=}}\max_{\pi_{k}(S_{k})}\Biggl{\{}\mathrm{E}_{W_{k}}\Biggl{[}r\bigl{(}\tilde{S}_{k},\pi_{k}(S_{k})\bigr{)}+\mathrm{E}_{S_{k+1}}\Bigl{[}\mathrm{E}_{W_{k+1}}\Bigl{[}\tilde{v}_{k+1}^{*}$
@@
=$\stackrel{{\scriptstyle(e)}}{{=}}\max_{\pi_{k}(S_{k})}\Biggl{\{}\mathrm{E}_{W_{k}}\Biggl{[}r\bigl{(}\tilde{S}_{k},\pi_{k}(S_{k})\bigr{)}+\mathrm{E}_{S_{k+1}}\Bigl{[}\tilde{v}_{k+1}^{*}(S_{k+1})\Bigr{]}$
@@
=$\stackrel{{\scriptstyle(f)}}{{=}}\max_{\pi_{k}(S_{k})}\Biggl{\{}\mathrm{E}_{W_{k}}\Biggl{[}r\bigl{(}\tilde{S}_{k},\pi_{k}(S_{k})\bigr{)}+\mathrm{E}_{S_{k+1}}\Bigl{[}\tilde{v}_{k+1}^{*}(S_{k+1})\Bigr{]}$
@@
=$\stackrel{{\scriptstyle(g)}}{{=}}\max_{\pi_{k}(S_{k})}\Bigl{\{}\mathrm{E}_{W_{k}}\bigl{[}r\bigl{(}\tilde{S}_{k},\pi_{k}(S_{k})\bigr{)}+\tilde{v}_{k+1}^{*}(S_{k+1})|S_{k},\pi_{k}(S_{k})\bigr{]}\Bigr{\}}$

### v3

@@ VI-B3 Method C
We estimate EVoMI and IVoMI for the vehicle-following control problem as given in Case 6. The velocity trajectory of the preceding vehicle i1i-1 is extracted from the Next Generation Simulation (NGSIM) dataset [40] when training the following vehicle. The optimal policy $\tilde{\pi}_{i}^{*}$ for the augmented-state SSDP in Case 4 is learned by the Twin Delayed Deep Deterministic policy gradient (TD3) algorithm [41]. This policy is also used as the acting policy $\pi_{i}$ when the predecessor’s acceleration is not available. In this case, $acc_{i-1,k}$ is set to 0 in the state.
The EVoMI is 3.12-3.12, which is calculated by evaluating the policy over $100$ test episodes in which $acc_{i-1,k}$ is either consistently unavailable or consistently available throughout the entire episode, and then computing the difference in the average total reward between the two settings.
@@ VII-B3 Across Multiple Control Intervals - When and What to Communicate
The “When to Communicate” and “What to Communicate” decisions can also be made statically only once at control interval 0 or semi-statically for every $Z>1$ control intervals. In this case, the subset of information or processed information will always be transmitted at every control interval within the effective time range of the decision. An example is choosing the information topology for a platoon [35].
@@
We consider a typical urban C-V2X network, where V2V links coexist with V2I links. A V2I link connects a vehicle to the BS and can be used for high-throughput services. A V2V link connects a pair of preceding vehicle i1i-1 and following vehicle $i$ for periodic transmission of cooperative awareness messages (CAM) to support advanced driving services. Specifically, at the beginning of each control interval $k$, the acceleration $acc_{i-1,k}$ of the predecessor i1i-1 is sampled to form the CAM and buffered in a queue before transmission to the corresponding follower $i$.
@@
At each communication interval $(k,t)$, the predecessor i1i-1 transmits the CAM generated at the beginning of each control interval $k$. Any CAM that is not fully transmitted by the end of the last communication interval (k,T1)(k,T-1) of control interval $k$ will be discarded. In other words, any CAM not fully delivered during the previous control interval k1k-1 is replaced by newly generated CAM at the beginning of control interval $k$.
@@
As introduced in Case 1, the control agent in follower $i$ determines its control input based on the local sensory information $ep_{i,k}$, $ev_{i,k}$, $acc_{i,k}$, as well as the remote sensory information received from its predecessor i1i-1.
@@
There is a communication agent in each predecessor i1i-1 to make communication decisions, which fall into one or more of the following categories:
@@
  • Every Communication Interval - How to Communicate: The communication agent makes RRM decisions at each communication interval $(k,t)$. To enhance spectrum utilization, we consider a setting where $L$ V2V links share the uplink resources orthogonally allocated to $M$ V2I links. With a slight abuse of notation, each V2V link, connecting a predecessor i1i-1 and its corresponding follower $i$, is indexed by $i-1\in\mathcal{L}=\{1,3,5,\ldots,2L-1\}$. Without loss of generality, we assume that every V2I link $m\in\mathcal{M}=\{0,\cdots,M-1\}$ is pre-assigned sub-channel $m$ and transmits with a constant power $P^{\rm I}_{m}$. Moreover, one or more V2V links can reuse the sub-channels of the V2I links for CAM transmission. We use the binary allocation indicator $\theta_{i-1,m,(k,t)}\in\{0,1\}$ to indicate whether V2V link i1i-1 occupies sub-channel $m$ at communication interval $(k,t)$ or not. Moreover, we consider that each V2V link i1i-1 occupies at most one sub-channel, i.e., $\sum_{m=0}^{M-1}\theta_{i-1,m,(k,t)}\leq 1$. Let $P^{\rm V}_{i-1,m,(k,t)}$ denote the transmit power of V2V link i1i-1 over the sub-channel $m$ at communication interval $(k,t)$. Therefore, each communication agent i1i-1 has to make the subchannel allocation $\{\theta_{i-1,m,(k,t)}\}_{m\in\mathcal{M}}$ and power allocation $\{P^{\rm V}_{i-1,m,(k,t)}\}_{m\in\mathcal{M}}$ decisions at each communication interval $(k,t)$.
@@
  • Every Control Interval - When to Communicate: Let $\phi_{i-1,k}\in\{0,1\}$ be the binary transmission indicator to indicate whether the sampled CAM $acc_{i-1,k}$ will be transmitted to the follower $i$ or not. At each control interval $k$, each communication agent i1i-1 can make decisions on $\phi_{i-1,k}$.
@@
  • Across Multiple Control Intervals - When to Communicate: Each communication agent i1i-1 makes a one-time decision on whether it will transmit its acceleration or not to its follower $i$. Once the decision is made, the predecessor will always transmit $acc_{i-1,k}$ or not at each control interval $k$.
@@ VII-C Optimization Objective
The instantaneous channel gain of V2V link i1i-1 over sub-channel $m$ (occupied by V2I link $m$) at communication interval $(k,t)$ is denoted by $G_{i-1,m,(k,t)}$. $G_{i-1,m,(k,t)}$ remains constant within a communication interval, and can be written as
@@
Similarly, let $G_{m,(k,t)}$ denote the instantaneous channel gain of V2I link $m$; $G_{i-1,B,m,(k,t)}$ the interference channel gain from V2V link i1i-1 transmitter at predecessor i1i-1 to V2I link $m$ receiver at the BS; $G_{B,i-1,m,(k,t)}$ the interference channel gain from V2I link $m$ transmitter at a vehicle to V2V link i1i-1 receiver at follower $i$; and $G_{j-1,i-1,m,(k,t)}$ the interference channel gain from V2V link j1j-1 transmitter at predecessor j1j-1 to V2V link i1i-1 receiver at follower $i$ over the sub-channel $m$.
The SINR $\gamma_{m,(k,t)}$ of V2I link $m$ and the SINR $\gamma_{i-1,m,(k,t)}$ of V2V link i1i-1 on sub-channel $m$ at communication interval $(k,t)$ are derived by
@@
respectively. $\sigma^{2}$ is the power of channel noise which satisfies the independent Gaussian distribution with a zero mean value. $I_{i-1,m,(k,t)}$ in (38) is the total interference power received by V2V link i1i-1 over sub-channel $m$, where
@@
[…]
The instantaneous data rates $C_{m,(k,t)}$ and $C_{i-1,(k,t)}$ of V2I link $m$ and V2V link i1i-1 at communication interval $(k,t)$ are respectively derived as
$C_{m,(k,t)}=B\log_{2}(1+\gamma_{m,(k,t)}),$ (39)
@@
Let $C^{\rm CAM}_{i-1,(k,t)}$ denote the transmission rate of V2V link i1i-1 in terms of CAM at communication interval $(k,t)$, which is given by
$C^{\rm CAM}_{i-1,(k,t)}=\frac{C_{i-1,(k,t)}}{N_{c}},$ (41)
@@
Each predecessor $i-1\in\mathcal{L}$ maintains a buffer to store its CAMs. Let $q^{\rm CAM}_{i-1,(k,t)}$ denote the queue length, i.e., the number of CAMs in the buffer of predecessor i1i-1 at communication interval $(k,t)$. At the beginning of the first communication interval $(k,0)$ of each control interval $k\in\mathcal{K}$, predecessor i1i-1 samples the driving status for control interval $k$ and generates a CAM. If the communication agent decides to transmit this CAM ($\phi_{i-1,k}=1$), it is stored in the buffer. During each communication interval $(k,t)$, where $t\in\mathcal{T}$, the queue length $q^{\rm CAM}_{i-1,(k,t)}$ is decreased by $\Delta t\times C^{\rm CAM}_{i-1,(k,t)}$, which represents the number of CAMs transmitted during this communication interval.
Since any CAM not fully transmitted by the end of the last communication interval (k,T1)(k,T-1) is discarded, the queue is reset at the start of each new control interval $k+1$. Specifically, at communication interval $(k+1,0)$, $q^{\rm CAM}_{i-1,(k+1,0)}$ is set to $1$ if a CAM is scheduled for transmission ($\phi_{i-1,k+1}=1$), or 0 otherwise ($\phi_{i-1,k+1}=0$). Because the queue length changes abruptly at the boundaries between control intervals, we use $q^{\rm CAM}_{i-1,(k,T)}$ to denote the queue length at the end of control interval $k$, i.e., after the final communication interval (k,T1)(k,T-1). Note that $q^{\rm CAM}_{i-1,(k,T)}\neq q^{\rm CAM}_{i,(k+1,0)}$ due to this discontinuity.
@@
$\left\{\begin{array}[]{ll}\phi_{i-1,k},&\mathrm{if}\quad t=0\\ \max\left[0,q^{\rm CAM}_{i-1,(k,t-1)}-C^{\rm CAM}_{i-1,(k,t-1)}\Delta t\right],&\mathrm{if}\quad 0<t\leq T\\ \end{array}\right.$. (44)
@@
With option (2), when the most recent available CAM of follower $i$ at the beginning of control interval $k$ corresponds to the CAM sampled at the predecessor i1i-1 in control interval $k-\tau_{i,k}$, we refer to $\tau_{i,k}$ as the observation delay at V2V receiver $i$, which is derived from the queue length as
@@
$J^{\mathrm{CM}}=\kappa_{1}\sum_{m=0}^{M-1}\mathrm{E}_{\pi^{\rm CM}}\left[\sum_{k=0}^{K-1}\sum_{t=0}^{T-1}\gamma^{kT+t}C_{m,(k,t)}\right]+\kappa_{2}\sum_{i-1\in\mathcal{L}}\Xi_{i},$ (48)
@@ VII-D VoI-oriented Optimization for Vehicular Communication
    $r^{\rm CM}_{k}=\kappa_{1}\left(\sum_{t=0}^{T-1}\gamma^{t}r_{(k,t)}^{\mathrm{CM\_trad}}\right)+\kappa_{2}\xi_{k+1},$ (50)
@@
Note that the $r_{(k,t)}^{\mathrm{CM\_trad}}$ component is present in the reward of every communication interval, while the IVoII component, i.e., $\xi_{k+1}$, is only present in the last communication interval of each control interval when t=T1t=T-1. The imperfect observation received by the control agent at control interval $k+1$ is affected by the degree of communication nonideality the transmitted data has experienced, which is in part a result of the sequence of decisions on communication actions $a_{(k,0)}^{\text{CM}},a_{(k,1)}^{\text{CM}},\ldots,a_{(k,T-1)}^{\text{CM}}$ that are selected during control interval $k$ by the communication agent.
@@
      $r^{\rm CM}_{(k,t)}$ $=\kappa_{1}\left(\sum_{m=0}^{M-1}\sum_{t=0}^{T-1}\gamma^{t}C_{m,(k,t)}\right)$
@@
      $\left\{\begin{array}[]{ll}\kappa_{1}\sum_{m=0}^{M-1}C_{m,(k,t)},&\ 0\leq t<T-1\\ \kappa_{1}\sum_{m=0}^{M-1}C_{m,(k,t)}+&\\ \kappa_{2}\sum_{i-1\in\mathcal{L}}\xi_{i,k+1},&\ t=T-1\\ \end{array}\right.$. (58)
@@ VIII Conclusion and Future Work
$\mathrm{Pr}(f^{S}(S_{k},a_{k},f^{W}(S_{k},a_{k},\zeta_{k}))|S_{k},a_{k})$
$=$ $\mathrm{Pr}(f^{S}(S_{k},a_{k},f^{W}(S_{k},a_{k},\zeta_{k}))|S_{k},a_{k},\{S_{k^{\prime}},a_{k^{\prime}}\}_{k^{\prime}\leq k-1}),$ (60)
which follows from the system’s state transition function $S_{k+1}=f^{S}(S_{k},a_{k},W_{k})$ given in Definition 1. In addition, $W_{k}=f^{W}(\tilde{W}_{k})$ according to the exogenous information function in Definition 1, while $\tilde{W}_{k}\subseteq\{S_{k},a_{k},\zeta_{k}\}$ as stated in Theorem 1.
@@
$\mathrm{Pr}(\zeta_{k}|S_{k},a_{k})$
$=$ $\mathrm{Pr}(\zeta_{k}|S_{k},a_{k},\{S_{k^{\prime}},a_{k^{\prime}}\}_{k^{\prime}\leq k-1})$
$=$ $\mathrm{Pr}(\zeta_{k}|S_{k},a_{k},\{S_{k^{\prime}},a_{k^{\prime}},W_{k^{\prime}}\}_{k^{\prime}\leq k-1}),$
@@
Since $W_{k}=f^{W}(S_{k},a_{k},\zeta_{k})$, $\zeta_{k}$ is a set of random variables whose distribution is independent of $S_{k}$ and $a_{k}$ by the expression of the formula. Thus, from (-A), we have
$\mathrm{Pr}(\zeta_{k})=\mathrm{Pr}(\zeta_{k}|\{S_{k^{\prime}},a_{k^{\prime}},W_{k^{\prime}}\}_{k^{\prime}\leq k-1}),$ (62)
which shows that the distribution of $\zeta_{k}$ does not depend on the past states, actions and exogenous information, i.e., $\{S_{k^{\prime}},a_{k^{\prime}},W_{k^{\prime}}\}_{k^{\prime}\leq k-1}$.
@@
$=\begin{pmatrix}f^{S}(S_{k},a_{k},W_{k})\\ f^{W}(f^{S}(S_{k},a_{k},W_{k}),\tilde{\pi}({f^{S}(S_{k},a_{k},W_{k})}),\zeta_{k})\end{pmatrix}$
$=f^{\tilde{S}}(\tilde{S}_{k},a_{k},\zeta_{k}),$ (65)
where the exogenous information $\zeta_{k}$ is an independent random variable with given distribution.
@@
$\stackrel{{\scriptstyle(b)}}{{=}}\max_{\pi_{k}(S_{k})}\Biggl{\{}\mathrm{E}_{W_{k}}\Biggl{[}r\bigl{(}\tilde{S}_{k},\pi_{k}(S_{k})\bigr{)}+\mathrm{E}_{W_{k+1}}\Bigl{[}\tilde{v}_{k+1}^{*}$
@@
$\stackrel{{\scriptstyle(c)}}{{=}}\max_{\pi_{k}(S_{k})}\Biggl{\{}\mathrm{E}_{W_{k}}\Biggl{[}r\bigl{(}\tilde{S}_{k},\pi_{k}(S_{k})\bigr{)}+\mathrm{E}_{S_{k+1}}\Bigl{[}\mathrm{E}_{W_{k+1}}\Bigl{[}\tilde{v}_{k+1}^{*}$
@@
$\stackrel{{\scriptstyle(d)}}{{=}}\max_{\pi_{k}(S_{k})}\Biggl{\{}\mathrm{E}_{W_{k}}\Biggl{[}r\bigl{(}\tilde{S}_{k},\pi_{k}(S_{k})\bigr{)}+\mathrm{E}_{S_{k+1}}\Bigl{[}\mathrm{E}_{W_{k+1}}\Bigl{[}\tilde{v}_{k+1}^{*}$
@@
$\stackrel{{\scriptstyle(e)}}{{=}}\max_{\pi_{k}(S_{k})}\Biggl{\{}\mathrm{E}_{W_{k}}\Biggl{[}r\bigl{(}\tilde{S}_{k},\pi_{k}(S_{k})\bigr{)}+\mathrm{E}_{S_{k+1}}\Bigl{[}\tilde{v}_{k+1}^{*}(S_{k+1})\Bigr{]}$
@@
$\stackrel{{\scriptstyle(f)}}{{=}}\max_{\pi_{k}(S_{k})}\Biggl{\{}\mathrm{E}_{W_{k}}\Biggl{[}r\bigl{(}\tilde{S}_{k},\pi_{k}(S_{k})\bigr{)}+\mathrm{E}_{S_{k+1}}\Bigl{[}\tilde{v}_{k+1}^{*}(S_{k+1})\Bigr{]}$
@@
$\stackrel{{\scriptstyle(g)}}{{=}}\max_{\pi_{k}(S_{k})}\Bigl{\{}\mathrm{E}_{W_{k}}\bigl{[}r\bigl{(}\tilde{S}_{k},\pi_{k}(S_{k})\bigr{)}+\tilde{v}_{k+1}^{*}(S_{k+1})|S_{k},\pi_{k}(S_{k})\bigr{]}\Bigr{\}}$